OpenStudy (anonymous):

f(x)=x^2 - x + 5/4 Graph by hand I can't use my graphing calculator.

OpenStudy (triciaal):

one approach set up a table with the values of x and the corresponding values for f(x) then plot the points another way find the value of f(0) and the value when f(x) = 0 this will give you three points the minimum needed to get a sketch for a parabola

OpenStudy (triciaal):

when x = 0 y = 5/4 when f(x) = 0 then x^2 - x + 5/4 = 0 solve for x you have 3 points do the sketch